OpenStudy (anonymous):

Use the sum-to-product formula to find the exact value: If \cos 255^\circ - \cos 195^\circ = \frac{\sqrt{A}}{2}

OpenStudy (anonymous):

\[\cos 255^\circ - \cos 195^\circ = \frac{\sqrt{A}}{2}\]

OpenStudy (anonymous):

The exact value of what?

OpenStudy (anonymous):

of A

OpenStudy (anonymous):

cos A - cos B= -2 sin(A+B)/2 * sin (A-B)/2

OpenStudy (anonymous):

Can you find it now?

OpenStudy (anonymous):

im working on it now

OpenStudy (anonymous):

the thing is should I make A the subject in that equation and solve it?

OpenStudy (anonymous):

Did you find cos 255- cos 195 already?

OpenStudy (anonymous):

-2sin225*sin15

OpenStudy (anonymous):

=.366

OpenStudy (anonymous):

what do i do next? I tried solving -2sin225*sin15 and equating it to \[\sqrt{A}/2\] and solving for A but it was wrong